OpenStudy (algtrigcalc):

A rock contains 0.9375 mg of uranium the day it was found and it was formed 4.2*10^9 years ago, how much uranium was present in the rock the day it was formed? the half life of uranium is 7*10^8 years.

OpenStudy (algtrigcalc):

how do i solve it using the formula ln(final/initial) =-kt k= decay constant

OpenStudy (aaronq):

It's more convenient to rearrange the formula for what you're solving for, then plug in your values. \(\sf\huge [A]_i=\dfrac{[A]_f}{e^{-kt}}\), where \(\sf k=\dfrac{ln(2)}{t_{1/2}}\) \(\sf\huge [A]_i=\dfrac{ 0.9375 ~mg}{e^{-(\frac{ln(2)}{7*10^8 years}* 4.2*10^9 years)} }\) \(\sf \huge [A]_i=59.999~mg \approx 60~ mg\)